# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 6.2 Reinforcement Learning-Based Methods
Reinforcement learning (RL) has emerged as a pivotal approach for aligning LLMs with complex, goal-driven tasks in scientific discovery. In traditional reinforcement learning, a Markov Decision Process (MDP) serves as the mathematical framework for representing the key components of an RL algorithm: states, actions, transition probabilities, rewards, and the discount factor. In the context of language models’ post-training, these elements are analogous to the input prompt plus the tokens generated so far (states), selecting the next token (action), transition probabilities conditioned on state and action , and the accuracy of the generated response for a given query (reward). RL has been widely adopted to enhance the reasoning capabilities of LLMs within this framework (deepseekai2025deepseekr1incentivizingreasoningcapability)(wang2023math)(shao2024deepseekmathpushinglimitsmathematical). Traditional RL frameworks have primarily focused on optimizing terminal rewards, such as achieving the correct solution or maximizing accuracy. However, scientific discovery and hypothesis generation are inherently process-oriented: intermediate steps such as problem decomposition, hypothesizing alternatives, self-correction, experiment design for hypothesis validation and citing credible sources are just as critical as the final output. To address this, an increasing body of research is shifting toward PRMs (lightman2023letsverifystepstep), which evaluate not only "what" models say but also "how" they think. ORMs, though easier to implement, tend to favor shallow heuristics and are susceptible to reward hacking. In contrast, PRMs provide granular, intermediate feedback that fosters deeper chains of reasoning. This evolution in RL design—from black-box evaluation to interpretable, step-wise reward shaping—enables new paradigms like debate, self-play, and simulation-based learning, all of which are increasingly relevant for equipping LLMs with the capacities for rigorous and transparent scientific reasoning.
@@
Table 2: Parameter-Adaptation Methods for Scientific Idea Generation (SIG) and Discovery (SD)
[…]
@@
CycleResear-cher (weng2025cycleresearcher) Scientific Idea Generation, Automated Research Iterative preference training (SimPO-based RL between policy and reward model) • Automates hypothesis–review cycles • Generates ideas close to accepted paper quality. • Limited cross-domain generalization • Reward unstability
@@
Learning to Generate Research Ideas with Dynamic Control (li2024learninggenerateresearchidea) Scientific Idea Generation, Creativity Control SFT $\rightarrow$ PPO-based controllable RL using multi-objective rewards (novelty, feasibility, usefulness) • Dynamic adjustment of generation with dimensional controllers • Balances novelty, feasibility, and effectiveness • Reward hacking risk • Trade-offs between novelty and feasibility
@@
CRPO (2025crpo) General Creative Generation Preference-based RL (modular DPO with creativity metrics: novelty, diversity, surprise, quality; trained on MUCE dataset) • Enhances originality and diversity while preserving quality • Outperforms DPO, DDPO, and GPT-4o on creativity benchmarks • LLM-as-a-judge can be exploited • Static preference datasets may bias learning
@@
Algorithm Discovery with LLMs (surina2025algorithmdiscoveryllmsevolutionary) Computational Scientific Discovery, Algorithmic Innovation Hybrid evolutionary search with RL fine-tuning (DPO + forward KL regularization) • Discovers novel algorithms for TSP, Bin Packing, symbolic regression • Improves optimality gaps and solution diversity • High computational cost • Search space explosion • Limited interpretability
@@
DARWIN Series (xie2023darwinseriesdomainspecific) Domain-Specific Scientific Discovery, Scientific Reasoning Domain SFT on $\sim$60K scientific tasks $\rightarrow$ multi-task fine-tuning with synthetic instruction generation • Improves domain reasoning and factual correctness • Strong performance on SciQ and related benchmarks • Domain overfitting • Risk of learning spurious features • Dependence on data diversity/quality
@@
SmileyLlama (cavanagh2025smileyllamamodifyinglargelanguage) Domain-Specific Discovery, Molecular Generation SFT $\rightarrow$ DPO $\rightarrow$ RL (iMiner algorithm) for property-optimized molecule generation • Generates chemically valid and diverse molecules • Achieves higher novelty and diversity than RNN baselines • Dataset bias and limited chemical coverage • Validation requires costly simulations
@@
ChemLM (kallergis2025chemlm) Domain-Specific Scientific Discovery, Chemistry Self-supervised molecular pretraining $\rightarrow$ Domain adaptation $\rightarrow$ SFT for property prediction • Improves molecular property prediction • Supports discovery of promising chemical candidates • Limited interpretability • Generalization across chemical families remains difficult
@@
ChemMLLM (tan2025chemmllmchemicalmultimodallarge) Domain-Specific Discovery, Multimodal Chemistry Multimodal fine-tuning across text, SMILES, and molecular images; unified encoder–decoder architecture • Bridges chemical modalities for visual–text–structure reasoning • Enables molecule generation and structure–function mapping • Complex modality alignment • High compute cost • Bias toward patterns seen in pretraining
@@ 6.2.1 Algorithm Discovery and Reasoning
RL has been instrumental in enabling LLMs to discover efficient algorithms and enhance scientific reasoning capabilities. For instance, surina2025algorithmdiscoveryllmsevolutionary demonstrated that RL-based fine-tuning, integrated within evolutionary search processes, dynamically improves LLM performance in algorithm discovery tasks. By using FunSearch (FunSearch) (only evolutionary search without RL) as a baseline, they showed that the combination of RL finetuning with evolutionary search can enhance the performance of the model and surpass the baseline in constructing optimal algorithms for tasks like bin packing, traveling salesman problem and flatpack. Additionally, leveraging expert domain knowledge, RL-based fine-tuning has been successfully applied to genomics, where expert forum discussions are automatically transformed into RL-friendly multiple-choice questions (MCQs), significantly enhancing the model’s scientific reasoning abilities (yin2025scientificreasoningllmstraining).
@@ 6.2.2 Domain-Specific Reasoning and Optimization
In protein discovery, RL has been coupled with Protein Language Models (pLMs) to design novel biologically plausible sequences optimized for therapeutic objectives. For instance, stocco2025guidinggenerativeproteinlanguage demonstrate how PPO-based fine-tuning can produce protein sequences with improved binding affinity and stability, including candidates for EGFR binders. Similarly, subramanian2024reinforcementlearningsequencedesign applies policy-gradient training on top of pre-trained protein models to generate diverse sequences tailored to objectives such as structural integrity. These studies illustrate how combining RL with pLMs enables protein discovery beyond natural sequence space.
In drug discovery, RL leverages structured and composite rewards to optimize molecular candidates. DrugImproverGPT introduces Structured Policy Optimization (SPO) to align molecules with therapeutic objectives while preserving beneficial properties (liu2025drugimprovergptlargelanguagemodel). Similarly, PPO-based fine-tuning balances protein-target efficacy and chemical validity, enhancing novelty and drug-likeness (ahmed2025improvingtargetedmoleculegeneration).
@@ 6.3 Hybrid and Preference Optimization Methods
One prominent application is research idea generation, where a two-stage framework combines SFT with controllable RL to dynamically balance novelty, feasibility, and effectiveness through multi-dimensional reward modeling (li2024learninggenerateresearchidea). Similarly, in molecular generation, a two-stage pipeline of SFT followed by RL significantly enhances structural diversity in generated molecule sets, outperforming conventional decoding methods (jang2025llmsgeneratediversemolecules).
@@
Preference optimization has emerged as a particularly effective complement to these hybrid strategies. DPO (rafailov2023dpo) simplifies alignment by directly learning from human preference pairs, avoiding the instability of complex reward models. SmileyLlama (cavanagh2025smileyllamamodifyinglargelanguage), derived from the Llama-3.1-8B-Instruct model, employs a pipeline of SFT combined with DPO to generate drug-like molecules optimized for biological activity and validated efficacy against protein targets. However, standard DPO often favors “safe” high-reward responses, limiting creativity and diversity in outputs.
@@
To address this, several recent works propose creativity- and diversity-enhanced preference optimization. CrPO introduces a multi-dimensional alignment framework that injects signals of novelty, diversity, surprise, and quality into the optimization process, producing outputs that are not only accurate but also more surprising and original (2025crpo). DivPO extends this idea by explicitly selecting preference pairs to reward rare but high-quality responses while penalizing common low-quality ones, leading to substantial improvements in output diversity without sacrificing quality (2025divpo). These approaches are particularly relevant for scientific idea generation, where generating multiple distinct but plausible hypotheses is as important as generating feasible ones.
@@
A recent example of this direction is CycleResearcher, which proposes a complete framework for automating the research cycle with open-source LLMs (weng2025cycleresearcher). The framework integrates two models: CycleResearcher, responsible for literature review, idea development, manuscript generation, and refinement; and CycleReviewer, which simulates peer review and provides iterative feedback via RL. The models are trained on two new datasets—Review-5k, containing structured peer reviews, and Research-14k, which provides outlines and segmented academic papers for fine-tuning. Through an iterative preference optimization approach (SimPO), the CycleResearcher–CycleReviewer loop allows the policy model to improve over multiple iterations, with CycleReviewer acting as a reward model to guide preference-based updates. Unlike standard DPO, which tends to reward only “safe” responses, this iterative design promotes more nuanced alignment with academic standards. In practice, the framework achieved competitive review scores, showing that LLMs can produce outputs near the quality of human preprints while still leaving room for further improvement.
@@
Structured optimization methods further extend this landscape by employing more specialized alignment techniques. Energy Rank Alignment (ERA) (chennakesavalu2025aligningtransformerscontinuousfeedback) uses gradient-based optimization to align molecular and protein generation with explicit chemical objectives, balancing robustness, diversity, and property alignment. Contrastive Preference Optimization (CPO) (gkoumas2024almolalignedlanguagemoleculetranslation) improves precision by explicitly contrasting optimal and suboptimal outputs, thereby enhancing reliability and mitigating issues like memorization or hallucination.
@@ 6.4 Takeaways and Limitations
SFT instills reliable reasoning patterns and aligns outputs with human demonstrations, ensuring consistency and interpretable results on tasks close to the training distribution. However, it is constrained by static datasets and can overfit to dataset biases, limiting generalization, handling of ambiguous prompts, and exploration of novel or high-risk hypotheses (chu2025sftmemorizesrlgeneralizes; hua2025intuitivefinetuningsimplifyingalignment).
RL guides exploration toward task-specific objectives, enabling creativity and risk-taking for potentially groundbreaking hypotheses. Its challenges include reward over-optimization (model focuses on easily achievable rewards, limiting novelty), reward hacking (model exploits loopholes in the reward function to get high reward without truly solving the task), output instability (early training can produce incoherent or repetitive outputs), sensitivity to reward design (poorly specified rewards may penalize creative or cross-domain ideas), and computational cost (multiple sampling and updates make large-scale training expensive) for large-scale discovery (mohammadi2024creativity; 2023arXiv231006452K; deepseekai2025deepseekr1incentivizingreasoningcapability).
@@
In summary, SFT ensures consistency, RL drives exploration, and hybrid methods balance both. While these approaches face challenges such as dataset biases, reward mis-specification, overfitting, and limited generalization, they internalize creativity and domain knowledge within the model itself. This reduces reliance on test-time computation or external resources and shifts the source of creativity from purely Process or Press signals toward the model’s internal capabilities, aligning with the Person aspect of 4p framework (rhodes1961analysis). Addressing the limitations while leveraging these intrinsic strengths could be the key to unlocking LLMs’ full potential for scientific hypothesis generation and discovery.

### v2

@@ 6.2 Reinforcement Learning-Based Methods
Reinforcement learning (RL) has emerged as a pivotal approach for aligning LLMs with complex, goal-driven tasks in scientific discovery. In traditional reinforcement learning, a Markov Decision Process (MDP) serves as the mathematical framework for representing the key components of an RL algorithm: states, actions, transition probabilities, rewards, and the discount factor. In the context of language models’ post-training, these elements are analogous to the input prompt plus the tokens generated so far (states), selecting the next token (action), transition probabilities conditioned on state and action , and the accuracy of the generated response for a given query (reward). RL has been widely adopted to enhance the reasoning capabilities of LLMs within this framework (DeepSeek-AI et al., 2025)(Wang et al., 2023b)(Shao et al., 2024). Traditional RL frameworks have primarily focused on optimizing terminal rewards, such as achieving the correct solution or maximizing accuracy. However, scientific discovery and hypothesis generation are inherently process-oriented: intermediate steps such as problem decomposition, hypothesizing alternatives, self-correction, experiment design for hypothesis validation and citing credible sources are just as critical as the final output. To address this, an increasing body of research is shifting toward PRMs (Lightman et al., 2023), which evaluate not only "what" models say but also "how" they think. ORMs, though easier to implement, tend to favor shallow heuristics and are susceptible to reward hacking. In contrast, PRMs provide granular, intermediate feedback that fosters deeper chains of reasoning. This evolution in RL design—from black-box evaluation to interpretable, step-wise reward shaping—enables new paradigms like debate, self-play, and simulation-based learning, all of which are increasingly relevant for equipping LLMs with the capacities for rigorous and transparent scientific reasoning.
@@
Table 2: Parameter-adaptation methods for scientific idea generation and discovery.
Method Field Adaptation strategy Main contribution / result Main challenges
@@
Cycle-Researcher (Weng et al., 2025a) Automated scientific research Iterative preference training (SimPO-based RL between policy and reward) • Automates hypothesis–review cycles • Produces ideas near accepted paper quality • Limited cross-domain generalization • Reward instability
@@
Dynamic Control (Li et al., 2024c) Scientific idea generation SFT $\rightarrow$ PPO-based controllable RL with multi-objective rewards (novelty, feasibility, usefulness) • Dynamically adjusts idea generation • Balances novelty and feasibility • Reward hacking risk • Trade-offs between objectives
@@
CRPO (Ismayilzada et al., 2025a) General creative generation Preference-based RL (modular DPO) using creativity metrics (novelty, diversity, surprise, quality) on MUCE dataset • Boosts originality and diversity with quality retention • Outperforms DPO, DDPO, GPT-4o • LLM-judge vulnerability • Static preference bias
@@
Algorithm Discovery (Surina et al., 2025) Comput-ational scientific discovery Hybrid evolutionary search + RL fine-tuning (DPO + forward KL) • Discovers novel algorithms for TSP, Bin Packing, regression • Improves optimality and diversity • High computation cost • Large search space • Limited interpretability
@@
DARWIN Series (Xie et al., 2023) Scientific reasoning Domain SFT on $\sim$60K tasks $\rightarrow$ multi-task fine-tuning with synthetic instruction generation • Improves domain reasoning and factual acc • Strong on SciQ benchmark • Domain overfitting • Spurious learning • Data diversity dependence
@@
Smiley-Llama (Cavanagh et al., 2025) Domain-specific discovery, molecular SFT $\rightarrow$ DPO $\rightarrow$ RL (iMiner) for property-optimized molecule design • Generates valid, diverse molecules • Higher novelty and diversity than RNN baselines • Dataset bias, limited coverage • Costly validation
@@
Chem-LM (Kallergis et al., 2025) Domain-specific discovery, chemistry Self-supervised molecular pretraining $\rightarrow$ domain adaptation $\rightarrow$ SFT for property prediction • Improves molecular property prediction • Aids discovery of chemical candidates • Limited interpretability • Poor cross-family generalization
@@
Chem-MLLM (Tan et al., 2025) Domain-specific discovery, chemistry Multimodal fine-tuning on text, SMILES, and molecular images; unified encoder–decoder design • Bridges modalities for visual–text–structure reasoning • Enables molecule generation / mapping • Complex alignment • High compute demand • Pretraining bias
@@ 6.2.1 Algorithm Discovery and Reasoning
RL has been instrumental in enabling LLMs to discover efficient algorithms and enhance scientific reasoning capabilities. For instance, Surina et al. (2025) demonstrated that RL-based fine-tuning, integrated within evolutionary search processes, dynamically improves LLM performance in algorithm discovery tasks. By using FunSearch (Romera-Paredes et al., 2024) (only evolutionary search without RL) as a baseline, they showed that the combination of RL finetuning with evolutionary search can enhance the performance of the model and surpass the baseline in constructing optimal algorithms for tasks like bin packing, traveling salesman problem and flatpack. Additionally, leveraging expert domain knowledge, RL-based fine-tuning has been successfully applied to genomics, where expert forum discussions are automatically transformed into RL-friendly multiple-choice questions (MCQs), significantly enhancing the model’s scientific reasoning abilities (Yin et al., 2025b).
@@ 6.2.2 Domain-Specific Reasoning and Optimization
In protein discovery, RL has been coupled with Protein Language Models (pLMs) to design novel biologically plausible sequences optimized for therapeutic objectives. For instance, Stocco et al. (2025) demonstrate how PPO-based fine-tuning can produce protein sequences with improved binding affinity and stability, including candidates for EGFR binders. Similarly, Subramanian et al. (2024) applies policy-gradient training on top of pre-trained protein models to generate diverse sequences tailored to objectives such as structural integrity. These studies illustrate how combining RL with pLMs enables protein discovery beyond natural sequence space.
In drug discovery, RL leverages structured and composite rewards to optimize molecular candidates. DrugImproverGPT introduces Structured Policy Optimization (SPO) to align molecules with therapeutic objectives while preserving beneficial properties (Liu et al., 2025d). Similarly, PPO-based fine-tuning balances protein-target efficacy and chemical validity, enhancing novelty and drug-likeness (Ahmed and Mohammed, 2025).
@@ 6.3 Hybrid and Preference Optimization Methods
One prominent application is research idea generation, where a two-stage framework combines SFT with controllable RL to dynamically balance novelty, feasibility, and effectiveness through multi-dimensional reward modeling (Li et al., 2024c). Similarly, in molecular generation, a two-stage pipeline of SFT followed by RL significantly enhances structural diversity in generated molecule sets, outperforming conventional decoding methods (Jang et al., 2025).
@@
Preference optimization has emerged as a particularly effective complement to these hybrid strategies. DPO (Rafailov and others, 2023) simplifies alignment by directly learning from human preference pairs, avoiding the instability of complex reward models. SmileyLlama (Cavanagh et al., 2025), derived from the Llama-3.1-8B-Instruct model, employs a pipeline of SFT combined with DPO to generate drug-like molecules optimized for biological activity and validated efficacy against protein targets. However, standard DPO often favors “safe” high-reward responses, limiting creativity and diversity in outputs.
@@
To address this, several recent works propose creativity- and diversity-enhanced preference optimization. CrPO introduces a multi-dimensional alignment framework that injects signals of novelty, diversity, surprise, and quality into the optimization process, producing outputs that are not only accurate but also more surprising and original (Ismayilzada et al., 2025a). DivPO extends this idea by explicitly selecting preference pairs to reward rare but high-quality responses while penalizing common low-quality ones, leading to substantial improvements in output diversity without sacrificing quality (Lanchantin et al., 2025). These approaches are particularly relevant for scientific idea generation, where generating multiple distinct but plausible hypotheses is as important as generating feasible ones.
@@
A recent example of this direction is CycleResearcher, which proposes a complete framework for automating the research cycle with open-source LLMs (Weng et al., 2025a). The framework integrates two models: CycleResearcher, responsible for literature review, idea development, manuscript generation, and refinement; and CycleReviewer, which simulates peer review and provides iterative feedback via RL. The models are trained on two new datasets—Review-5k, containing structured peer reviews, and Research-14k, which provides outlines and segmented academic papers for fine-tuning. Through an iterative preference optimization approach (SimPO), the CycleResearcher–CycleReviewer loop allows the policy model to improve over multiple iterations, with CycleReviewer acting as a reward model to guide preference-based updates. Unlike standard DPO, which tends to reward only “safe” responses, this iterative design promotes more nuanced alignment with academic standards. In practice, the framework achieved competitive review scores, showing that LLMs can produce outputs near the quality of human preprints while still leaving room for further improvement.
@@
Structured optimization methods further extend this landscape by employing more specialized alignment techniques. Energy Rank Alignment (ERA) (Chennakesavalu et al., 2025) uses gradient-based optimization to align molecular and protein generation with explicit chemical objectives, balancing robustness, diversity, and property alignment. Contrastive Preference Optimization (CPO) (Gkoumas, 2024) improves precision by explicitly contrasting optimal and suboptimal outputs, thereby enhancing reliability and mitigating issues like memorization or hallucination.
@@ 6.4 Takeaways and Limitations
SFT instills reliable reasoning patterns and aligns outputs with human demonstrations, ensuring consistency and interpretable results on tasks close to the training distribution. However, it is constrained by static datasets and can overfit to dataset biases, limiting generalization, handling of ambiguous prompts, and exploration of novel or high-risk hypotheses (Chu et al., 2025; Hua et al., 2025a).
RL guides exploration toward task-specific objectives, enabling creativity and risk-taking for potentially groundbreaking hypotheses. Its challenges include reward over-optimization (model focuses on easily achievable rewards, limiting novelty), reward hacking (model exploits loopholes in the reward function to get high reward without truly solving the task), output instability (early training can produce incoherent or repetitive outputs), sensitivity to reward design (poorly specified rewards may penalize creative or cross-domain ideas), and computational cost (multiple sampling and updates make large-scale training expensive) for large-scale discovery (Mohammadi, 2024; Kirk et al., 2023; DeepSeek-AI et al., 2025).
@@
Empirical evidence supports this characterization. In the CycleResearcher (Weng et al., 2025a) study, parameter-adapted models consistently outperform the AI Scientist baseline in peer-review–style evaluation, achieving higher overall scores (5.15–5.38 vs. 4.31) and non-zero acceptance rates (24–35% vs. 0%). At a criterion level, adapted models show improvements in soundness, presentation, and contribution, with the latter—interpreted as a proxy for novelty—improving from 2.15 to 2.53–2.60 approaching the performance of human-reviewed preprints across all three dimensions. Complementary evidence from the CrPO (Ismayilzada et al., 2025a) study, evaluated using NoveltyBench (Zhang et al., 2025c), further clarifies while SFT substantially improves quality-interpreted as valueness- ($\approx$ 5.0 vs. 3.1–4.1 for base LLMs vs. 4.3–4.9 for CrPO) but yields more limited gains in novelty ($\approx$ 5.7 vs. 4.0 for base LLMs), remaining well below preference based approaches (e.g., CrPO: 6.8–7.9). These results indicate that parameter adaptation primarily strengthens valueness, with stronger novelty emerging when optimization explicitly rewards exploration. Further comparisons can be found in Table 3.
In summary, SFT ensures consistency, RL drives exploration, and hybrid methods balance both. While these approaches face challenges such as dataset biases, reward mis-specification, overfitting, and limited generalization, they internalize creativity and domain knowledge within the model itself. This reduces reliance on test-time computation or external resources and shifts the source of creativity from purely Process or Press signals toward the model’s internal capabilities, aligning with the Person aspect of 4p framework (Rhodes, 1961). Addressing the limitations while leveraging these intrinsic strengths could be the key to unlocking LLMs’ full potential for scientific hypothesis generation and discovery.
Table 3: Method categories and their expected effects on novelty and valueness in scientific ideation. Expected effects reflect empirically observed trends and theoretically motivated interpretations reported in prior work, rather than guaranteed outcomes.
Method Category Primary Mechanism (4Ps) Expected Effect on Novelty Expected Effect on Valueness Representative Works
Knowledge & Retrieval Augmentation External knowledge injection (Press) $\uparrow$$\leftrightarrow$ Combinatorial novelty via recombination of retrieved concepts; may overfit to retrieved context $\uparrow$ Grounding, correctness, and feasibility; reduced hallucinations Section 2.3 (Gu et al., 2025; Gao et al., 2025b)
Prompt-based Distributional Steering Input-level guidance (Press / Process) $\uparrow$ Novelty by steering generation toward lower-probability regions (constraints, role-play, reframing) $\uparrow$$\downarrow$ Structured prompting can improve reasoning; misaligned constraints may degrade value Section 3.5 (Zhao et al., 2025; Li et al., 2024a)
Inference-time Scaling & Search Systematic exploration beyond single-pass decoding (Process) $\uparrow\uparrow$ Exploratory novelty via branching, refinement, and structured search $\uparrow$$\downarrow$ Improves with evaluation or pruning; unconstrained search risks invalid ideas Section 4.4 (Sprueill et al., 2023; Feng et al., 2025)
Multi-agent & Deliberative Systems Interaction, debate, and critique (Process; proxy for Person) $\uparrow\uparrow\uparrow$ Potential for transformational novelty via emergent interactions $\uparrow$ Improved coherence, plausibility and even robustness through iterative critique Section 5.3 (Gottweis et al., 2025; Johnson and Davis, 2024)
Parameter Adaptation & Learning Internalization of knowledge and strategies (Person) $\leftrightarrow$$\uparrow$ SFT, which reinforces training patterns; RL or PO via internalized exploration $\uparrow$ Consistency, feasibility, and robustness through internalized constraints Section 6.4 (Weng et al., 2025a; Ismayilzada et al., 2025a)